# On the multiplier spectrum of polynomials

### v1

@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
$C_{\infty}(F)\cap{\rm PGL}_{2}(\mathbb{C})={\rm Aut}_{\infty}(F)=E_{0}(F)=E(F)\cap{\rm PGL}_{2}(\mathbb{C}).$
To show $C_{\infty}(F)=E(F)$, it suffices to prove $E(F)\setminus E_{0}(F)\subseteq C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})$. Let $g\in E(F)\setminus E_{0}(F)$. Then $\deg(g)\geq 2$ and $\mu_{g}=\mu_{F}$. By [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ g^{\circ l}$. Clearly, $k_{1}>k_{2}$. Applying Theorem 3.6 recursively, we get $F^{\circ(k_{1}-k_{2})}=g^{\circ l}$. Then $g$ commutes with $F^{\circ(k_{1}-k_{2})}=g^{\circ l}$, so $g\in C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})$. Hence, $C_{\infty}(F)=E(F)$.
Clearly, $\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\subseteq C_{\infty}(F)$. For maps of degree $1$,
$\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\cap{\rm PGL}_{2}(\mathbb{C})={\rm Aut}_{\infty}(F)=C_{\infty}(F)\cap{\rm PGL}_{2}(\mathbb{C}).$
To show $\left\langle{\rm Aut}_{\infty}(F),F\right\rangle=C_{\infty}(F)$, it suffices to prove
$C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})\subseteq\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\setminus{\rm PGL}_{2}(\mathbb{C}).$
Let $G\in C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})$. By [Rit23], there exist integers $k,l>0$ such that $F^{\circ k}=G^{\circ l}$. By Corollary 3.9, there exists $\mu_{1}\in{\rm PGL}_{2}(\mathbb{C})$ such that $G=\mu_{1}^{-1}\circ F^{\circ s}$, where $s=k/l\in\mathbb{Z}_{>0}$ (take $\mu_{1}(z)=z$ and $s=k$ when $l=1$). Then
$F^{\circ sl}=F^{\circ k}=G^{\circ l}=(\mu_{1}^{-1}\circ F^{\circ s})^{\circ l}.$
By [Pak25a, Lemma 3.3], $\mu_{1}^{-1}\in{\rm Aut}(F^{\circ sl})\subseteq{\rm Aut}_{\infty}(F)$. Then $G=\mu_{1}^{-1}\circ F^{\circ s}\in\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\setminus{\rm PGL}_{2}(\mathbb{C})$.
We have shown that $C_{\infty}(F)=\left\langle{\rm Aut}_{\infty}(F),F\right\rangle=E(F)$ holds. ∎
We now consider general polynomials. Here, “general” means the property holds for polynomials in a non-empty Zariski open subset of the parameter space ${\rm Poly}^{m}(\mathbb{C})$.
@@
Up to conjugation by linear polynomials, we only need to consider monic centered polynomials of degree $m$, i.e., polynomials of the form
@@
where $a_{m-2},\dots,a_{0}\in\mathbb{C}$. The parameter space of monic centered polynomials of degree $m$ is
${\rm Poly}_{mc}^{m}=\{(a_{0},\dots,a_{m-2})\}\cong\mathbb{A}^{m-1}.$
The moduli space ${\rm MPoly}^{m}$ is the quotient of ${\rm Poly}_{mc}^{m}$ by the finite cyclic group $\mathbb{U}_{m-1}$ of (m1)(m-1)-th roots of unity acting diagonally on $\mathbb{A}^{m-1}\cong{\rm Poly}_{mc}^{m}$ by
@@
See [FG22, §2.1] for more details. Let $R(t)$ be the resultant ${\rm Res}_{z}(F^{\prime}(z),F(z)-t)$ in $z$. Then
$R(t)=m^{m}\prod_{\zeta:F^{\prime}(\zeta)=0}(F(\zeta)-t)\in(\mathbb{C}(a_{0},\dots,a_{m-2}))^{{\rm alg}}[t].$
Set $Z:={\rm Res}_{t}(R(t),R^{\prime}(t))\in\mathbb{C}[a_{0},\dots,a_{m-2}]\setminus\{0\}$ (the existence of pre-simple $f\in{\rm Poly}_{mc}^{m}(\mathbb{C})$ implies $Z\neq 0$). The resultant $Z$ defines a hypersurface in ${\rm Poly}_{mc}^{m}\cong\mathbb{A}^{m-1}$, corresponding to polynomials that are not pre-simple. Hence, a general polynomial of degree $m$ is pre-simple. ∎
@@
Embed ${\rm Rat}_{m}$ into $\mathbb{P}^{2m+1}$ via the coefficients of rational maps (see [Sil12]). By the proof of [Pak25a, Lemma 3.7], there exists a closed subvariety $Z$ of $\mathbb{P}^{2m+1}$ over $\mathbb{C}$ such that
$\{F\in{\rm Rat}_{m}(\mathbb{C}):G(F)\neq 1\}=Z\cap{\rm Rat}_{m}(\mathbb{C})$
and $F\notin Z$ for some monic centered polynomial $F\in{\rm Poly}_{mc}^{m}(\mathbb{C})$ of degree $m$. The conclusion follows. ∎
Lemmas 3.15, 3.16, and Theorem 3.14 immediately imply:
@@
For every integer $m\geq 4$, a general polynomial $F\in\mathbb{C}[z]$ of degree $m$ satisfies $E_{0}(F)={\rm Aut}_{\infty}(F)=G_{0}(F)=1$ and $E(F)=C_{\infty}(F)=\left\langle F\right\rangle$.
@@
$\mathcal{A}:=\{z\in\mathbb{C}:F^{\circ r}(z)\in c^{*}(\mathcal{O}),\deg_{z}(F)=\cdots=\deg_{F^{\circ(r-1)}(z)}(F)=1\}.$
Since $F$ is pre-simple, induction shows $\mathcal{A}\geq(m-2)^{r}k\geq 2k$. As $f:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map, we have $\mathcal{A}\subseteq c^{*}(\mathcal{O})$. Then $2k\leq\#\mathcal{A}\leq\#c^{*}(\mathcal{O})=k$, which is a contradiction. Hence, $F^{\circ r}$ is not a generalized Lattès map for any $r\in\mathbb{Z}_{>0}$. ∎
@@
According to [Pak23, §2.3], for a rational map $f\in\mathbb{C}(z)$ of degree $\geq 2$ not of monomial type, $f$ is a generalized Lattès map if and only if $f^{\circ r}$ is a generalized Lattès map for some $r\in\mathbb{Z}_{>0}$, which is equivalent to $f^{\circ r}$ being a generalized Lattès map for all $r\in\mathbb{Z}_{>0}$. Thus, by Lemma 3.12, it suffices to prove Lemma 3.18 for $r=1$.
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$, $G\in\mathbb{C}[z]$ a polynomial of degree $\geq 4$, and $X\in\mathbb{C}[z]\setminus\mathbb{C}$ a non-constant polynomial such that $X\circ G=F^{\circ r}\circ X$ for some $r\in\mathbb{Z}_{>0}$. Then $l:=\log_{m}(\deg(X))$ is a non-negative integer, and there exists $\nu\in{\rm PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ l}\circ\nu$ and $G=\nu^{-1}\circ F^{\circ r}\circ\nu$. If $G$ and $X$ are polynomials, we may take $\nu$ to be a linear polynomial.
@@
If $\deg(X)=1$, the result is trivial. Assume $\deg(X)\geq 2$. By Lemma 3.18, $F^{\circ r}$ is not a generalized Lattès map. Applying [Pak25a, Theorem 4.1], there exist $Y\in\mathbb{C}(z)\setminus\mathbb{C}$ and $d\in\mathbb{Z}_{>0}$ such that $X\circ Y=F^{\circ rd}$. By Corollary 3.9, $l:=\log_{m}(\deg(X))$ is a positive integer, and there exists $\mu\in{\rm PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ l}\circ\mu$. Then
@@
By Theorem 3.6, we get $F^{\circ r}\circ\mu=\mu\circ G$, so $G=\mu^{-1}\circ F^{\circ r}\circ\mu$. If $G$ and $X$ are polynomials, then $Y$ is also a polynomial, and by Corollary 3.9, we may take $\mu$ to be a linear polynomial. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then for every $k\in\mathbb{Z}_{>0}$, the map $\gamma:{\rm Aut}(F^{\circ k})\to{\rm Aut}(F^{\circ k})$ (as in Lemma 3.10) is a group automorphism.
@@
By Theorem 3.14, ${\rm Aut}_{\infty}(F)=G_{0}(F)$. Fix $k\in\mathbb{Z}_{>0}$. Then ${\rm Aut}(F^{\circ k})\subseteq{\rm Aut}_{\infty}(F)=G_{0}(F)$. For every $\nu\in{\rm Aut}(F^{\circ k})$, we have
@@
$=F\circ\nu\circ F^{\circ k}=\gamma(\nu)\circ F\circ F^{\circ k}=\gamma(\nu)\circ F^{\circ k}\circ F,$
so $F^{\circ k}\circ\gamma(\nu)=\gamma(\nu)\circ F^{\circ k}$. Hence, $\gamma(\nu)\in{\rm Aut}(F^{\circ k})$. By Lemma 3.10, the map $\gamma:{\rm Aut}(F^{\circ k})\to{\rm Aut}(F^{\circ k})$ is an injective group endomorphism on the finite group ${\rm Aut}(F^{\circ k})$, hence an automorphism. ∎
We now describe periodic curves for endomorphisms $(F_{1},F_{2})$ on $(\mathbb{P}^{1})^{2}$, where $F_{1},F_{2}\in\mathbb{C}[z]$ are pre-simple of the same degree $\geq 4$.
@@
Let $F_{1},F_{2}\in\mathbb{C}[z]$ be pre-simple polynomials of the same degree $m\geq 4$, and let $C\subset(\mathbb{P}^{1})^{2}$ be an irreducible algebraic curve over $\mathbb{C}$ that is not a vertical or horizontal line. Then for every $k\in\mathbb{Z}_{>0}$, the following are equivalent:
@@
    There exist $s\in\mathbb{Z}_{\geq 0}$, $\alpha\in{\rm PGL}_{2}(\mathbb{C})$, and $\nu\in{\rm Aut}(F_{1}^{\circ k})$ such that F_2^k=αF_1^kα^-1, and $C$ is one of the graphs y=(ανF_1^s)(x) or x=(νF_1^sα^-1)(y), where $x$ and $y$ are the coordinates on $(\mathbb{P}^{1})^{2}$.
@@
Assume (i) holds for some $k\in\mathbb{Z}_{>0}$. By Lemma 3.18, $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are not generalized Lattès maps. By [Pak25a, Theorem 4.2], there exist non-constant rational maps $X_{1},X_{2},Y_{1},Y_{2},B\in\mathbb{C}(z)\setminus\mathbb{C}$ and $\tilde{k}\in\mathbb{Z}_{\geq 0}$ such that:
@@
By Theorem 3.20, there exist $\beta,\alpha\in{\rm PGL}_{2}(\mathbb{C})$ and $d_{1},d_{2}\in\mathbb{Z}_{\geq 0}$ such that
@@
Conjugating $B$ suitably, we may assume $\beta(z)=z$. Then $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),(F_{2}^{\circ d_{2}}\circ\alpha)(t))$ and $F_{1}^{\circ k}=B=\alpha^{-1}\circ F_{2}^{\circ k}\circ\alpha$. Since
@@
we have $\alpha^{-1}\circ F_{2}\circ\alpha\in E(F_{1})=\left\langle{\rm Aut}_{\infty}(F_{1}),F_{1}\right\rangle$ by Theorem 3.14. Also, ${\rm Aut}_{\infty}(F_{1})=G_{0}(F)$ by Theorem 3.14. Hence, $\alpha^{-1}\circ F_{2}\circ\alpha=\tau\circ F_{1}$ for some $\tau\in{\rm Aut}_{\infty}(F_{1})=G_{0}(F)$. Since $F_{1}^{\circ k}=(\alpha^{-1}\circ F_{2}\circ\alpha)^{\circ k}=(\tau\circ F_{1})^{\circ k}$, we get $\tau\in{\rm Aut}(F_{1}^{\circ k})$ by [Pak25a, Lemma 3.3]. Note that $F_{2}=\alpha\circ\tau\circ F_{1}\circ\alpha^{-1}$ and $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),\alpha\circ(\tau\circ F_{1})^{\circ d_{2}}(t))$. By Lemma 3.21, $(\tau\circ F_{1})^{\circ d_{2}}=\tau^{\prime}\circ F_{1}^{\circ d_{2}}$ for some $\tau^{\prime}\in{\rm Aut}(F_{1}^{\circ k})$. Then $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),(\alpha\circ\tau^{\prime}\circ F_{1}^{\circ d_{2}}(t))$.
If $d_{1}\leq d_{2}$, then $C$ is parametrized by $t\mapsto(t,(\alpha\circ\tau^{\prime}\circ F_{1}^{\circ(d_{2}-d_{1})}(t))$, and we are done by setting $\nu:=\tau^{\prime}$ and $s:=d_{2}-d_{1}$.
If $d_{1}>d_{2}$, then by Lemma 3.21,
@@
for some $\tau^{\prime\prime}\in{\rm Aut}(F_{1}^{\circ k})$. Then $C$ is parametrized by
@@
Following [FG22], two polynomials $f,g\in\mathbb{C}[z]$ of degree $\geq 2$ are intertwined if there exists an algebraic curve $Z\subset(\mathbb{A}^{1})^{2}$ (possibly reducible) whose projections to both axes are onto, and $Z$ is invariant under the endomorphism $(f,g):(\mathbb{A}^{1})^{2}\to(\mathbb{A}^{1})^{2}$. The following theorem is a polynomial version of [JX25, Theorem 3.3]:
Theorem 3.23 (Theorem 1.5).
For every integer $d\geq 4$, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for every $f\in U$:
@@
    for every pre-simple $g\in{\rm Poly}^{d}(\mathbb{C})$, if $f$ and $g$ are intertwined, then $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.

### v2

@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
$C_{\infty}(F)\cap\mathrm{PGL}_{2}(\mathbb{C})=\mathrm{Aut}_{\infty}(F)=E_{0}(F)=E(F)\cap\mathrm{PGL}_{2}(\mathbb{C}).$
To show $C_{\infty}(F)=E(F)$, it suffices to prove $E(F)\setminus E_{0}(F)\subseteq C_{\infty}(F)\setminus\mathrm{PGL}_{2}(\mathbb{C})$. Let $g\in E(F)\setminus E_{0}(F)$. Then $\deg(g)\geq 2$ and $\mu_{g}=\mu_{F}$. By [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ g^{\circ l}$. Clearly, $k_{1}>k_{2}$. Applying Theorem 3.7 recursively, we get $F^{\circ(k_{1}-k_{2})}=g^{\circ l}$. Then $g$ commutes with $F^{\circ(k_{1}-k_{2})}=g^{\circ l}$, so $g\in C_{\infty}(F)\setminus\mathrm{PGL}_{2}(\mathbb{C})$. Hence, $C_{\infty}(F)=E(F)$.
Clearly, $\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle\subseteq C_{\infty}(F)$. For Möbius transformations,
$\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle\cap\mathrm{PGL}_{2}(\mathbb{C})=\mathrm{Aut}_{\infty}(F)=C_{\infty}(F)\cap\mathrm{PGL}_{2}(\mathbb{C}).$
To show $\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle=C_{\infty}(F)$, it suffices to prove
$C_{\infty}(F)\setminus\mathrm{PGL}_{2}(\mathbb{C})\subseteq\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle\setminus\mathrm{PGL}_{2}(\mathbb{C}).$
Let $G\in C_{\infty}(F)\setminus\mathrm{PGL}_{2}(\mathbb{C})$. By [Rit23] and Lemma 3.18, there exist integers $k,l>0$ such that $F^{\circ k}=G^{\circ l}$. By Corollary 3.12, there exists $\mu\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $G=\mu\circ F^{\circ s}$, where $s=k/l\in\mathbb{Z}_{>0}$ (take $\mu(z)=z$ when $l=1$). Then
$(F^{\circ s})^{\circ l}=F^{\circ k}=G^{\circ l}=(\mu\circ F^{\circ s})^{\circ l}.$
By [Pak25a, Lemma 3.5], $\mu\in\mathrm{Aut}((F^{\circ s})^{\circ l})\subseteq\mathrm{Aut}_{\infty}(F)$. Then
$G=\mu\circ F^{\circ s}\in\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle\setminus\mathrm{PGL}_{2}(\mathbb{C}).$
We obtain $\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle=C_{\infty}(F)$.
This establishes that $C_{\infty}(F)=E(F)=\left\langle\mathrm{Aut}_{\infty}(F),F\right\rangle\subseteq\mathbb{C}[z]$. ∎
For a fixed integer $m\geq 2$, we consider properties which are satisfied by general polynomials of degree $m\geq 2$. Here, “general” means that the property holds for all polynomials in a non-empty Zariski open subset of the parameter space $\mathrm{Poly}^{m}(\mathbb{C})$.
@@
Up to conjugacy, we only need to consider monic centered polynomials of degree $m$, i.e., polynomials of the form
@@
where $a_{m-2},\dots,a_{0}\in\mathbb{C}$. The parameter space of monic centered polynomials of degree $m$ is identified with
$\mathrm{Poly}_{mc}^{m}=\{(a_{0},\dots,a_{m-2})\}\cong\mathbb{A}^{m-1}.$
The moduli space $\mathrm{MPoly}^{m}$ is the quotient of $\mathrm{Poly}_{mc}^{m}$ by the finite cyclic group $\mathbb{U}_{m-1}$ of (m1)(m-1)-th roots of unity (in $\mathbb{C}$) acting diagonally on $\mathbb{A}^{m-1}\cong\mathrm{Poly}_{mc}^{m}$ by
@@
See [FG22, § 2.1] for more details. Let $R(t)$ be the resultant $\mathrm{Res}_{z}(F^{\prime}(z),F(z)-t)$ in $z$. Then
$R(t)=m^{m}\prod_{\zeta:F^{\prime}(\zeta)=0}(F(\zeta)-t)\in\mathbb{C}(a_{0},\dots,a_{m-2})[t].$
Set $Z:=\mathrm{Res}_{t}(R(t),R^{\prime}(t))\in\mathbb{C}[a_{0},\dots,a_{m-2}]$. Since there exist pre-simple polynomials in $\mathrm{Poly}_{mc}^{m}(\mathbb{C})$ (for example, $z^{m}-mz$ when $m\geq 3$, and $z^{2}-1$ when $m=2$), the polynomial $R(t)$ in $a_{0},\dots,a_{m-2}$ is not the zero polynomial. Then $Z$ defines the empty set or a hypersurface $H$ in $\mathrm{Poly}_{mc}^{m}\cong\mathbb{A}^{m-1}$, exactly corresponding to (monic centered) polynomials of degree $m$ that are not pre-simple. Hence, a general polynomial of degree $m$ is pre-simple. ∎
@@
Embed $\mathrm{Rat}_{m}$ into $\mathbb{P}^{2m+1}$ via the coefficients of rational maps of degree $m$ (see [Sil12]). By the proof of [Pak25a, Lemma 3.10], there exists a closed subvariety $Z$ of $\mathbb{P}^{2m+1}$ over $\mathbb{C}$ such that
$\{F\in\mathrm{Rat}_{m}(\mathbb{C}):G(F)\neq 1\}=Z\cap\mathrm{Rat}_{m}(\mathbb{C})\quad\text{and}\quad\mathrm{Poly}_{mc}^{m}(\mathbb{C})\setminus Z\neq\emptyset.$
The conclusion follows. ∎
Lemma 3.21, Lemma 3.22, and Theorem 3.20 immediately imply:
@@
For every integer $m\geq 4$, a general polynomial $F\in\mathbb{C}[z]$ of degree $m$ satisfies $G_{0}(F)=G(F)=\mathrm{Aut}_{\infty}(F)=E_{0}(F)=1$ and $C_{\infty}(F)=E(F)=\left\langle F\right\rangle$.
@@
$\mathcal{A}:=\{z\in\mathbb{C}\colon F^{\circ r}(z)\in c^{*}(\mathcal{O}),\deg_{z}(F)=\cdots=\deg_{F^{\circ(r-1)}(z)}(F)=1\}.$
Using the pre-simplicity of $F$, an induction on $r$ shows $\mathcal{A}\geq(m-2)^{r}k$. As $F^{\circ r}:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map, we have $\mathcal{A}\subseteq c^{*}(\mathcal{O})$. Then
$2k\leq(m-2)^{r}k\leq\#\mathcal{A}\leq\#c^{*}(\mathcal{O})=k,$
which is a contradiction. Hence, $F^{\circ r}$ is not a generalized Lattès map for any $r\in\mathbb{Z}_{>0}$. ∎
@@
According to [Pak23, § 2.3], the following conditions are equivalent for a rational map $f\in\mathbb{C}(z)$ of degree $\geq 2$:
  • $f$ is a generalized Lattès map;
  • $f^{\circ r}$ is a generalized Lattès map for some $r\in\mathbb{Z}_{>0}$;
  • $f^{\circ r}$ is a generalized Lattès map for all $r\in\mathbb{Z}_{>0}$.
Then it suffices to prove Lemma 3.24 for the case $r=1$.
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Let $G,X\in\mathbb{C}[z]\setminus\mathbb{C}$ be non-constant polynomials such that $X\circ G=F^{\circ r}\circ X$ for some $r\in\mathbb{Z}_{>0}$. Then $l:=\log_{m}(\deg(X))$ is a non-negative integer, and there exists $\mu\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ l}\circ\mu$ and $G=\mu^{-1}\circ F^{\circ r}\circ\mu$. If $G$ and $X$ are polynomials, we may take $\mu\in\mathrm{Aff}(\mathbb{C})$.
@@
If $\deg(X)=1$, then the result is trivial. Assume $\deg(X)\geq 2$. By Lemma 3.24, $F^{\circ r}$ is not a generalized Lattès map. Applying [Pak25a, Theorem 4.1], there exist $Y\in\mathbb{C}(z)\setminus\mathbb{C}$ and $d\in\mathbb{Z}_{>0}$ such that $X\circ Y=F^{\circ rd}$. By Corollary 3.12, $l:=\log_{m}(\deg(X))$ is a positive integer, and there exists $\mu\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ l}\circ\mu$. Then
@@
By Theorem 3.7, we get $F^{\circ r}\circ\mu=\mu\circ G$, so $G=\mu^{-1}\circ F^{\circ r}\circ\mu$. If $G$ and $X$ are polynomials, then $Y$ is also a polynomial, and we may take $\mu\in\mathrm{Aff}(\mathbb{C})$ by Corollary 3.12. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then for every $k\in\mathbb{Z}_{>0}$, the map $\gamma:\mathrm{Aut}(F^{\circ k})\to\mathrm{Aut}(F^{\circ k}),\sigma\mapsto\tau_{\sigma}$ (see (3.4)) is a well-defined group automorphism.
@@
By Theorem 3.20, $\mathrm{Aut}_{\infty}(F)=G_{0}(F)$. Fix $k\in\mathbb{Z}_{>0}$. Then $\mathrm{Aut}(F^{\circ k})\subseteq\mathrm{Aut}_{\infty}(F)=G_{0}(F)$; so the map $\gamma:\mathrm{Aut}(F^{\circ k})\to G_{0}(F)$ is well-defined. For every $\nu\in\mathrm{Aut}(F^{\circ k})$, we have
@@
$=F\circ\nu\circ F^{\circ k}=\gamma(\nu)\circ F\circ F^{\circ k}=\gamma(\nu)\circ F^{\circ k}\circ F;$
so $F^{\circ k}\circ\gamma(\nu)=\gamma(\nu)\circ F^{\circ k}$, i.e., $\gamma(\nu)\in\mathrm{Aut}(F^{\circ k})$. By Lemma 3.16, the map $\gamma:\mathrm{Aut}(F^{\circ k})\to\mathrm{Aut}(F^{\circ k})$ is an injective group endomorphism on the finite group $\mathrm{Aut}(F^{\circ k})$, hence an automorphism. ∎
We now describe periodic curves for endomorphisms $(F_{1},F_{2})$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$, where $F_{1},F_{2}\in\mathbb{C}[z]$ are pre-simple of the same degree $\geq 4$.
@@
Let $F_{1},F_{2}\in\mathbb{C}[z]$ be pre-simple polynomials of the same degree $d\geq 4$, and let $C\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$ be an irreducible algebraic curve over $\mathbb{C}$ that is neither a horizontal nor a vertical line. Then for every $k\in\mathbb{Z}_{>0}$, the following are equivalent:
@@
    There exist $s\in\mathbb{Z}_{\geq 0}$, $\alpha\in\mathrm{Aff}(\mathbb{C})$, and $\nu\in\mathrm{Aut}(F_{1}^{\circ k})$ such that F_2^k=αF_1^kα^-1, and $C$ is one of the graphs y=(ανF_1^s)(x) or x=(νF_1^sα^-1)(y), where $x$ and $y$ are the coordinates on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
@@
Let $k\in\mathbb{Z}_{>0}$ and assume (i) holds for $k$. By Lemma 3.24, $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are not generalized Lattès maps. By [Pak23, Theorem 1.1], there exist non-constant rational maps $X_{1},X_{2},Y_{1},Y_{2},B\in\mathbb{C}(z)\setminus\mathbb{C}$ and $\tilde{k}\in\mathbb{Z}_{\geq 0}$ such that:
@@
By Theorem 3.26, there exist $\beta,\alpha\in\mathrm{PGL}_{2}(\mathbb{C})$ and $d_{1},d_{2}\in\mathbb{Z}_{\geq 0}$ such that
@@
Conjugating $B$ suitably, we may assume $\beta(z)=z$. Then $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),(F_{2}^{\circ d_{2}}\circ\alpha)(t))$ and
$F_{1}^{\circ k}=B=\alpha^{-1}\circ F_{2}^{\circ k}\circ\alpha.$
For $1\leq i\leq 2$, since $F_{i}^{\circ k}$ is not a generalized Lattès map (in particular, not conjugate to $z^{d^{k}}$), $\infty$ is the unique totally invariant fixed point of $F_{i}^{\circ k}$. Hence the Möbius transformation $\alpha$ must fix the point $\infty$, so $\alpha\in\mathrm{Aff}(\mathbb{C})$.
Since
@@
we have $\alpha^{-1}\circ F_{2}\circ\alpha\in E(F_{1})=\left\langle\mathrm{Aut}_{\infty}(F_{1}),F_{1}\right\rangle$ by Theorem 3.20. Also, $\mathrm{Aut}_{\infty}(F_{1})=G_{0}(F)$ by Theorem 3.20. Hence, $\alpha^{-1}\circ F_{2}\circ\alpha=\tau\circ F_{1}$ for some $\tau\in\mathrm{Aut}_{\infty}(F_{1})=G_{0}(F)$. Since
$F_{1}^{\circ k}=(\alpha^{-1}\circ F_{2}\circ\alpha)^{\circ k}=(\tau\circ F_{1})^{\circ k},$
$\tau\in\mathrm{Aut}(F_{1}^{\circ k})$ by [Pak25a, Lemma 3.5]. Then $F_{2}=\alpha\circ\tau\circ F_{1}\circ\alpha^{-1}$ and $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),\alpha\circ(\tau\circ F_{1})^{\circ d_{2}}(t))$. By Lemma 3.27, $(\tau\circ F_{1})^{\circ d_{2}}=\tau^{\prime}\circ F_{1}^{\circ d_{2}}$ for some $\tau^{\prime}\in\mathrm{Aut}(F_{1}^{\circ k})$. Then $C$ is parametrized by
$t\mapsto(F_{1}^{\circ d_{1}}(t),(\alpha\circ\tau^{\prime}\circ F_{1}^{\circ d_{2}})(t)).$
If $d_{1}\leq d_{2}$, then $C$ is parametrized by
$t\mapsto(t,(\alpha\circ\tau^{\prime}\circ F_{1}^{\circ(d_{2}-d_{1})})(t)),$
and we are done by setting $\nu:=\tau^{\prime}$ and $s:=d_{2}-d_{1}$. If $d_{1}>d_{2}$, then by Lemma 3.27,
@@
for some $\tau^{\prime\prime}\in\mathrm{Aut}(F_{1}^{\circ k})$. Hence $C$ is parametrized by
@@
Recall that two polynomials $f,g\in\mathbb{C}[z]$ of degree $\geq 2$ are intertwined if there exists an algebraic curve $Z\subset\mathbb{A}^{2}$ (possibly reducible) whose projections to both axes are onto such that $Z$ is invariant under the endomorphism $(f,g):\mathbb{A}^{2}\to\mathbb{A}^{2}$. If $f$ and $g$ are intertwined, then $\deg(f)=\deg(g)$. The following theorem is a polynomial version of [JX25, Theorem 3.3]:
Theorem 3.29 (=Theorem 1.5).
For every integer $d\geq 4$, there exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for every $f\in U$:
@@
    for every pre-simple $g\in\mathrm{Poly}^{d}(\mathbb{C})$, $f$ and $g$ are intertwined if and only if $[f]=[g]$ in $\mathrm{MPoly}^{d}(\mathbb{C})$.